Assume inflation in Turkey is 50% and 2% in the US. Using the relative PPP, 

a) Calculate the exchange rate (𝑅 = 𝑇𝐿/$) and the exchange rate (𝑅 = $/𝑇𝐿).

b) Assume that the Turkish inflation increases to 100% while the US inflation remains at 2%, calculate the new exchange rate between the Turkish Lira and the US dollar.



1
Expert's answer
2022-02-28T11:34:25-0500

A) inflation in Turkey is 48% higher than in US

"R=\frac{TL}{\$}"

"R=\frac{1+0.5}{1+0.02}-1=0.4705"

"(1+0.4705)\times0.072=1.05876"


"R=\frac{\$}{TL}"

"R=\frac{1+0.02}{1+0.5}-1=-0.32"

"(1-0.32)\times13.82=9.3976"

b)"R=\frac{1+1.0}{1+0.02}-1=0.9608"

"(1+0.9608)\times0.072=1.412"